A SIDC JOCKY HAS 7 SONGS TO PLAY THREE ARE SLOW SONGS, AND FOUR ARE FAST SONGS WACH SONG IS TO BE PLAYED ONLY ONCE IN HOW MANY WAYS CAN THE DISC JOCKY PLAY THE 7 SONGS IF THE SONGS CAN BE PLAYED IN ANY ORDER; THE FIRST SONG MUST BE A SLOW SONG AND THE LAST SONG MUST BE A SLOW SONG; THE FIRST TWO SONGS MUST BE FAST SONGS
Answer by stanbon(57337)

Suppose the life of a certain brand of vacuum cleaner is normally distributed with a mean life of 12 years and a standard deviation of 4 years. A vacuum cleaner is randomly chosen.
10. For how many years should the manufacturer warranty its vacuum cleaners if it is willing to replace only 1% of the vacuum cleaners it sells (Round answer to the closest whole number).
-------------------
invNorm(0.01) = -2.3263
----
Use x = zs+u
--------------------------
x = -2.3263*4+12 = 2.6946 years = 3 yrs. to the closest whole number,
===================
Cheers,
Stan H.